-6x+5y2z=10 -4x+5y+5z=29 -9x+4y+z=8
Answer by RAY100(1637)

Assuming the 1st eqn is ...... 5y + 2z ....
,
1)-6x+5y+2z = 10,,,,,mult by 5....-30x+25y+10z=50
,
2)-4x +5y +5z =29,,,,mult by 2,,,,-8x +10y +10z =58
,
3)-9x +4y +z =8,,,,,mult by 10,,,,-90x +40y +10z = 80
,
This is three unknowns,,three eqns,,,therefore can be solved.
If you have a graphing calculator, and know augmented matrix, good spot to use, otherwise,
,
subtract (2) from (1) to get,,,-22x+15y = -8,,,,,,,mult by 2,,,-44x+30y+-16
substract (2) from(3) to get ,,,-82x+30y = 22
,
subtract again to get,,,-38x = 38,,,,x=(-1)ANSWER
,
substitute ,,,-22(-1) +15y = -8
22+15y = -8
15y = -30,,,,y=(-2) ANSWER
,
substitute in original (1),, -6(-1)+5(-2)+2z=10
6-10+2z = 10
2z = 14,,,,,,z=7 ANSWER
,
check by subst into original eqns
1) -6(-1)+5(-2)+2(7) = 10
6-10+14=10,,,,ok
,
2) -4(-1)+5(-2)+5(7) = 29
4-10+35 = 29,,,,,ok
,
3) -9(-1)+4(-2)+(7) =8
9-8+7 =8,,,,,,ok